# Spatial Phonons: A Phenomenological Viscous Dark Energy Model for DESI

### v2

@@ Abstract
1 Introduction
Over the last quarter century a coherent observational picture has emerged in which the expansion of the Universe is accelerating. The first direct evidence came from type Ia supernovae used as standardizable candles at redshifts of order unity, which appeared dimmer than expected in a decelerating Universe [18, 19]. These measurements showed that the cosmic scale factor has entered a late time phase of acceleration driven by a component with negative effective pressure that is now referred to as dark energy.
Subsequent probes have both confirmed this conclusion and measured the properties of the dark sector with increasing precision. Measurements of the cosmic microwave background anisotropies determined the geometry and matter density of the Universe and pointed to a spatially flat model in which roughly seventy percent of the present day energy density resides in a smooth dark energy component [47]. Large galaxy surveys revealed a characteristic baryon acoustic oscillation feature in the clustering pattern, which acts as a standard ruler and provides robust constraints on the expansion history through distance and Hubble rate measurements [27]. Combined analyses of supernovae, BAO and the CMB have established the current concordance picture in which cold dark matter and dark energy dominate the energy budget on cosmological scales [28, 36].
In most observational studies the dark energy sector is described phenomenologically by an equation of state parameter $w(z)\equiv p_{\rm DE}/\rho_{\rm DE}$ that may depend on redshift. The simplest possibility is a cosmological constant $\Lambda$ with strictly constant w=1w=-1. More general scenarios allow a mild time dependence, often represented by low dimensional parametrizations such as the Chevallier–Polarski–Linder form [20, 24],
$w(z)=w_{0}+w_{a}\frac{z}{1+z}$
which captures a wide class of smooth evolutions with only two parameters and is convenient for confronting different models with data [28, 36].
The apparent success of a cosmological constant as a fit to observations immediately raises the cosmological constant problem. The value of $\Lambda$ required to explain cosmic acceleration is extremely small when expressed in natural units, while naive estimates of the vacuum energy density from quantum field theory are larger by many orders of magnitude [13]. This enormous mismatch is difficult to explain within conventional effective field theory and has motivated both dynamical dark energy models and theories in which gravity is modified on large scales [44]. Reviews of the subject emphasize that understanding the origin and nature of dark energy is one of the central open problems in fundamental physics [28, 36].
On the observational side this challenge has led to a program of precision cosmology that combines multiple complementary probes. The Dark Energy Task Force report organized these efforts in terms of supernovae, baryon acoustic oscillations, weak gravitational lensing and galaxy clusters, and stressed the need for wide field spectroscopic surveys that can map the large scale structure over huge volumes and over a wide range of redshifts [29]. Baryon acoustic oscillation measurements play a central role in this program because they are based on linear physics at recombination and provide a robust standard ruler that is relatively insensitive to the complex astrophysics of galaxy formation [27].
@@ 1 Introduction
[…]
In this work we investigate the possibility that such a transient departure from w=1w=-1 is a manifestation of the viscoelastic response of space itself. We model space as a three dimensional brane with a uniform tension that plays the role of a bare cosmological constant and that can be identified with a Nambu–Goto type term in the gravitational action [8, 59]. On top of this background tension we introduce a phonon like medium on the brane, described within the effective field theory of continuous media by a triplet of scalar fields that label comoving volume elements [38, 40, 43]. The scalar invariant that measures compression determines both the bulk modulus and the longitudinal sound speed of these spatial phonons, while dissipation is encoded in an effective bulk viscosity that we model using a Maxwell type relaxation law inspired by Kubo formulas for bulk viscous stresses [7, 11, 16, 12, 48]. This structure produces an effective dark energy component whose equation of state $w_{\rm eff}(z)$ can temporarily dip below minus one during the epoch when the Hubble rate is comparable to a phonon gap scale and tends back toward a non phantom value at very early and very late times.
We identify a small set of dimensionless parameters that control this behaviour and show that for natural choices of these parameters the resulting $w_{\rm eff}(z)$ closely tracks the $w_{0}w_{a}$ form preferred by DESI based analyses [52, 51, 50]. The remainder of the paper develops the elastic brane and phonon fluid in detail, derives the corresponding cosmological background equations and effective bulk viscosity, and then compares the predicted evolution of $w_{\rm eff}(z)$ to the constraints from the first year DESI BAO data in combination with external probes.
@@ 2.1 Scalar Description of Longitudinal Spatial Phonons
In [59], it was proposed that physical space can be modelled as a fundamental elastic three brane. The brane is described at long distances by a Nambu–Goto [8, 17] type worldvolume action,
@@
where $T_{s}$ is the geometric tension of space and $g_{\mu\nu}$ is the spacetime metric. In the same work [59], the ‘no geometric sequester theorem’ was formulated. The theorem states that in any minimal matter vacuum sequester [41, 42] the global constraint cancels constant contributions that arise from the matter sector but leaves a purely geometric unit operator $\int d^{4}x\sqrt{-g}$ untouched. As a result the coefficient $T_{s}$ cannot be neutralized by the matter sequester and remains as a residual geometric contribution after renormalization from graviton loops and matter–graviton loops. This residual geometric tension can then be identified with the observed dark energy density at some scale $\mu_{\rm IR}$ appropriate for cosmology.
@@
In particular, we introduce three scalar fields [38, 40, 43],
@@ 2.2 Stress-Energy Tensor of the Phonon Fluid
The stress energy tensor is defined in the standard way by [25],
@@ 2.3 Calculating the Spatial Bulk Modulus
The adiabatic sound speed of longitudinal phonons is defined by (see [35] for fluids in FLRW and sound speed discussions),
@@ 3.1 Microscopic definition of bulk viscosity
In a relativistic fluid the bulk viscosity describes the dissipative response of the pressure to a homogeneous expansion or contraction. In covariant form the viscous part of the stress tensor can be written as (see [12, 9, 22, 33, 48, 39] for related discussion),
@@
The microscopic definition of $\zeta$ can be given through a Kubo formula in terms of the retarded correlator of the stress tensor [7, 11, 16]. A convenient expression is,
@@ 3.2 Physical picture on the Elastic Brane
In this work, we will interpret the Gibbons-Hawking temperature [10] of a spatially flat FRW spacetime,
@@
as an effective horizon temperature associated with the cosmological apparent horizon, rather than as literal Hawking radiation from an exact de Sitter event horizon. This viewpoint follows the horizon-thermodynamics approach in which the Friedmann equations can be written as a first law $dE=T_{H}dS+WdV$ at the apparent horizon, with $T_{H}=1/2\pi R_{A}$ and $R_{A}=1/H$ for $k=0$ FRW backgrounds [30, 31, 32, 37]. We assume that the phonon sector is in near-equilibrium with this effective thermal bath at temperature $T_{H}$ during the epoch when bulk viscosity is relevant. Concretely, we assume that the microscopic processes that create and annihilate phonons and that redistribute their momentum are fast enough that the phonon distribution relaxes toward a local equilibrium characterized by $T_{H}$ on time scales shorter than or comparable to the Hubble time during the epoch when bulk viscosity is relevant. Under this assumption the Kubo formula and the Maxwell viscoelastic model provide a meaningful characterization of the bulk viscosity $\zeta$ in terms of correlation functions evaluated at temperature $T_{H}$. We will discuss these ideas in the Sec. 3.3.
@@ 3.3 Maxwell Visco-Elastic Model and Cosmological Bulk Viscosity in FLRW
For a spatially flat FLRW metric [2, 3, 5, 6, 26], with scale factor $a(t)$ and comoving fluid, one has $\nabla_{\alpha}u^{\alpha}=3H$, where $H=\dot{a}/a$ is the Hubble parameter. The bulk viscous pressure $\Pi$ is then the spatial trace of equation (3.1), which gives,
@@
Note that a positive bulk viscosity produces a negative $\Pi$ during expansion [22, 33, 34, 15, 46].
A simple and physically transparent way to obtain $\zeta$ is provided by a Maxwell visco-elastic model for the bulk stress [1]. We assume that the bulk viscous pressure $\Pi$ obeys the simple linear causal relationship,
@@
We identify the combination $K_{\text{ph}}\tau(H)$ has the dimensions of viscosity and plays the role of the static bulk viscosity of the medium [21]. In a Maxwell model one typically has $\zeta_{0}(H)\sim K_{\text{ph}}\tau(H)$. This relation reflects the fact that the same microphysics that sets the elastic modulus also sets the relaxation time. Substituting this scaling relation gives,
@@
At temperature $T_{H}$, the interaction rate for processes that involve real phonon excitations typically behaves as [14],
@@
where $n_{\phi}$ is the phonon number density, $\sigma_{\rm ph}$ is an effective cross section and $v_{\rm rel}$ is a typical relative velocity. For a gapped excitation with mass $m_{\phi}$ at temperature $T_{H}$, the number density follows,
@@
For the purposes of this paper it is sufficient to treat $m_{\phi}$ and therefore $H_{\star}$ as phenomenological parameters and to require that the condition $T_{H_{\star}}\sim m_{\phi}$ is realized at some redshift $z_{\star}$ before the onset of dark energy domination. In essence, we assume that $H_{\star}$ is reached at some $z_{\star}>0.3$, in fact, we may assume that $H_{\star}$ is reached even before cosmic acceleration starts, that is $z_{\star}>0.6$. This ensures that the model produces a transient viscous epoch, with a mild phantom like deviation of the effective equation of state (see [23] for a phantom like equation of state for dark energy) followed by a rapid return to an effectively non viscous or at least not significantly viscous dark energy. In Sec. 3.4 next, we will discuss the effective equation of state with the phonon pressure contribution and highlighting the phantom dip scenario.
@@ 3.4 Effective Equation of State of Space and the Phantom Dip
4 Cosmological Implications and Comparison with DESI

### v3

@@ Abstract
We explore a phenomenological model of dark energy in which space is treated as an elastic brane with uniform tension $T_{s}$ and supports a longitudinal phonon sector described by three scalar fields $\phi^{I}$. At the background level the construction reproduces a perfect fluid whose enthalpy and bulk modulus are controlled by two dimensionless parameters $\varepsilon$ and $\kappa$, which encode the elastic and viscous response and determine both the effective equation of state $w_{\rm eff}(z)$ and the physical sound speed thriugh $c_{s}^{2}=\kappa/\varepsilon$. Motivated by the consistency requirements of a phonon effective description, we focus on the stable regime $0<\kappa<\varepsilon<1$, which prevents tachyonic behavior and is natural given that $\varepsilon$ and $\kappa$ are defined as fractional response coefficients. To connect the model predictions to current observational summaries, we map $w_{\rm eff}(z)$ onto the Chevallier–Polarski–Linder form $w(z)=w_{0}+w_{a}z/(1+z)$ over the redshift range relevant to DESI. We then use the public DESI DR1 compressed Gaussian likelihood in $(w_{0},w_{a})$ for BAO combined with Pantheon+ Type I a supernovae and Planck 2018, and scan the model parameter space to identify viable regions. We find that the model can reproduce the DESI compressed constraints at the background level, with representative best match values $(w_{0},w_{a})\simeq(-0.83,-0.74)$ and therefore, a near luminal sound speed $c_{s}^{2}\simeq 0.9618$ as an output of the fit rather than a built in restriction. The covariance weighted distance is found to satisfy $\chi^{2}\ll 1$ in the $(w_{0},w_{a})$ plane. The associated phonon mass scale $m_{\phi}=H_{\star}/(2\pi)$ is ultralight, corresponding to a Compton wavelength of order the cosmological horizon, consistent with interpreting the relevant excitations as infrared collective modes of the brane.
1 Introduction
Over the last quarter century a coherent observational picture has emerged in which the expansion of the Universe is accelerating. The first direct evidence came from type Ia supernovae used as standardizable candles at redshifts of order unity, which appeared dimmer than expected in a decelerating Universe (Riess et al. (1998); Perlmutter et al. (1999)). Subsequent measurements of the cosmic microwave background anisotropies and of baryon acoustic oscillations established a spatially flat cosmology in which a dominant smooth component with negative effective pressure drives the late time acceleration (Eisenstein et al. (2005); Copeland et al. (2006); Caldwell & Kamionkowski (2009); Aghanim et al. (2020)).
@@ 1 Introduction
$w(z)=w_{0}+w_{a}\,\frac{z}{1+z}\,$, (1.1)
which is flexible enough to represent a wide range of smooth evolutions while remaining directly constrainable by current datasets.
@@
Our goal is to provide a minimal and calculable mechanism in which a transient departure from w=1w=-1 arises as an effective macroscopic response of space itself. We model space as a three dimensional brane with a uniform background tension that plays the role of a bare cosmological constant and that admits a Nambu–Goto type description in the gravitational effective action (Goto (1971); Khan (2025)). On top of this background we introduce a phonon-like medium on the brane, described within the effective field theory of continuous media by a triplet of scalar fields that label comoving volume elements (Dubovsky et al. (2012); Endlich et al. (2013); Nicolis et al. (2014)).
In particular, Section 2 develops the elastic sector in a covariant and parameter-controlled way, identifying the small set of dimensionless combinations that govern the bulk modulus, longitudinal sound speed, and the background dark-energy-like response of the phonon medium. A central ingredient for obtaining a controlled transient phantom-like phase without introducing ghosts is dissipation and relaxation. Section 3 introduces a viscoelastic constitutive relation in which bulk viscous stress relaxes toward its instantaneous elastic value on a characteristic timescale, implemented via a Maxwell-type relaxation law motivated by the Kubo description of bulk viscous response (Kubo (1957); Hosoya et al. (1984); Jeon (1995); Landau & Lifshitz (1987); Bemfica et al. (2022)). When the Hubble rate becomes comparable to the phonon gap and relaxation scales, the effective pressure can temporarily become more negative, producing an epoch in which $w_{\rm eff}(z)$ dips below 1-1 before returning toward a non-phantom value at both early and late times.
Section 4 then connects this microphysical description to observations using a deliberately transparent compressed-likelihood strategy. Rather than re-running a full end-to-end likelihood analysis, we build a Gaussian approximation in the $(w_{0},w_{a})$ plane using the public DESI DR1 cosmology posterior products, specifically the Cobaya MCMC chains released for the DESI DR1 BAO cosmology analysis and its standard external dataset combinations (Aghanim et al. (2020); DESI Collaboration (2024); Scolnic et al. (2021); Brout et al. (2022); Torrado & Lewis (2021)). This yields a mean vector and covariance for $(w_{0},w_{a})$ that define a compact target for model comparison. In this sense our data interface is modular, easy to update when new DESI releases appear, and sufficient for establishing whether the viscoelastic phonon framework can naturally realize the type of transient CPL-like behaviour suggested by current combined analyses.
The remainder of the paper develops the background cosmology implied by the elastic and viscous sectors, maps the theory parameters to an effective $w_{\rm eff}(z)$ trajectory, and quantifies the agreement with the DESI-motivated benchmark within the compressed likelihood. We also summarize the physical interpretation of the key parameter combinations and outline how the framework can be extended to more complete perturbation-level tests and to future data releases.
@@ 2.1 Scalar Description of Longitudinal Spatial Phonons
In (Khan (2025)), it was proposed that physical space can be modelled as a fundamental elastic three brane. The brane is described by a Nambu–Goto (Goto (1971); Polchinski (1998)) type worldvolume action,
@@
where $T_{s}$ is the geometric tension of space and $g_{\mu\nu}$ is the spacetime metric. In the same work (Khan (2025)), the ‘no geometric sequester theorem’ was formulated. The theorem states that in any minimal matter vacuum sequester (Kaloper & Padilla (2014b, a)) the global constraint cancels constant contributions that arise from the matter sector but leaves a purely geometric unit operator $\int d^{4}x\sqrt{-g}$ untouched. As a result the coefficient $T_{s}$ cannot be neutralized by the matter sequester and remains as a residual geometric contribution after renormalization from graviton loops and matter–graviton loops. This residual geometric tension can then be identified with the observed dark energy density at some scale $\mu_{\rm IR}$ appropriate for cosmology.
@@
In particular, we introduce three scalar fields (Dubovsky et al. (2012); Endlich et al. (2013); Nicolis et al. (2014)),
@@ 2.2 Stress-Energy Tensor of the Phonon Fluid
The stress energy tensor is defined in the standard way by (Carroll (2004)),
@@ 2.3 Calculating the Spatial Bulk Modulus
The adiabatic sound speed of longitudinal phonons is defined by (see (Weinberg (2008)) for fluids in FLRW and sound speed discussions),
@@ 3.1 Microscopic definition of bulk viscosity
In a relativistic fluid the bulk viscosity describes the dissipative response of the pressure to a homogeneous expansion or contraction. In covariant form the viscous part of the stress tensor can be written as (see (Landau & Lifshitz (1987); Bemfica et al. (2022); Weinberg (1971); Zimdahl et al. (2001); Colistete et al. (2007); Kovtun (2012)) for related discussion),
@@
The microscopic definition of $\zeta$ can be given through a Kubo formula in terms of the retarded correlator of the stress tensor (Kubo (1957); Hosoya et al. (1984); Jeon (1995)). A convenient expression is,
@@ 3.2 Physical picture on the Elastic Brane
In this work, we will interpret the Gibbons-Hawking temperature (Gibbons & Hawking (1977)) of a spatially flat FRW spacetime,
@@
as an effective horizon temperature associated with the cosmological apparent horizon, rather than as literal Hawking radiation from an exact de Sitter event horizon. This viewpoint follows the horizon-thermodynamics approach in which the Friedmann equations can be written as a first law $dE=T_{H}dS+WdV$ at the apparent horizon, with $T_{H}=1/2\pi R_{A}$ and $R_{A}=1/H$ for $k=0$ FRW backgrounds (Akbar & Cai (2007); Cai & Cao (2007); Gong & Wang (2007); Sheykhi (2010)). We assume that the phonon sector is in near-equilibrium with this effective thermal bath at temperature $T_{H}$ during the epoch when bulk viscosity is relevant. Concretely, we assume that the microscopic processes that create and annihilate phonons and that redistribute their momentum are fast enough that the phonon distribution relaxes toward a local equilibrium characterized by $T_{H}$ on time scales shorter than or comparable to the Hubble time during the epoch when bulk viscosity is relevant. Under this assumption the Kubo formula and the Maxwell viscoelastic model provide a meaningful characterization of the bulk viscosity $\zeta$ in terms of correlation functions evaluated at temperature $T_{H}$. We will discuss these ideas in the Sec. 3.3.
@@ 3.3 Maxwell Visco-Elastic Model and Cosmological Bulk Viscosity in FLRW
For a spatially flat FLRW metric (Mukhanov (2005)), with scale factor $a(t)$ and comoving fluid, one has $\nabla_{\alpha}u^{\alpha}=3H$, where $H=\dot{a}/a$ is the Hubble parameter. The bulk viscous pressure $\Pi$ is then the spatial trace of equation (3.1), which gives,
@@
Note that a positive bulk viscosity produces a negative $\Pi$ during expansion (Zimdahl et al. (2001); Colistete et al. (2007); Singh (2008); Gariel & Le Denmat (1995); Czajka & Jeon (2019)).
A simple and physically transparent way to obtain $\zeta$ is provided by a Maxwell visco-elastic model for the bulk stress (Maxwell (1867)). We assume that the bulk viscous pressure $\Pi$ obeys the simple linear causal relationship,
@@
We identify the combination $K_{\text{ph}}\tau(H)$ has the dimensions of viscosity and plays the role of the static bulk viscosity of the medium (Roylance (2001)). In a Maxwell model one typically has $\zeta_{0}(H)\sim K_{\text{ph}}\tau(H)$. This relation reflects the fact that the same microphysics that sets the elastic modulus also sets the relaxation time. Substituting this scaling relation gives,
@@
At temperature $T_{H}$, the interaction rate for processes that involve real phonon excitations typically behaves as (Kolb & Turner (1990)),
@@
where $n_{\rm ph}$ is the phonon number density, $\sigma_{\rm ph}$ is an effective cross section and $v_{\rm ph}$ is a typical relative velocity. For a gapped excitation with mass $m_{\phi}$ at temperature $T_{H}$, the number density follows,
@@
For the purposes of this paper it is sufficient to treat $m_{\phi}$ and therefore $H_{\star}$ as phenomenological parameters and to require that the condition $T_{H_{\star}}\sim m_{\phi}$ is realized at some redshift $z_{\star}$ before the onset of dark energy domination. In essence, we assume that $H_{\star}$ is reached at some $z_{\star}>0.3$, in fact, we may assume that $H_{\star}$ is reached even before cosmic acceleration starts, that is $z_{\star}>0.6$. This ensures that the model produces a transient viscous epoch, with a mild phantom like deviation of the effective equation of state (see Caldwell (2002) for a phantom like equation of state for dark energy) followed by a rapid return to an effectively non viscous or at least not significantly viscous dark energy. In Sec. 3.4 next, we will discuss the effective equation of state with the phonon pressure contribution and highlighting the phantom dip scenario.
@@ 3.4 Effective Equation of State of Space and the Phantom Dip
4 Compressed Likelihood fit in the $(w_{0},w_{a})$ Plane
4.1 Scope and Motivation
In this section we compare the phonon model to late time expansion history constraints that are commonly summarized in the CPL plane $(w_{0},w_{a})$. The aim is to test whether the model can reproduce the region of the $(w_{0},w_{a})$ posterior preferred by current data, using a transparent and reproducible compressed likelihood.
We do not evaluate the BAO, supernova, or CMB likelihoods directly. Instead, we use publicly released MCMC chains for $w_{0}w_{a}$CDM and compress their joint posterior in $(w_{0},w_{a})$ to a bivariate Gaussian. The chains used here correspond to a combined analysis of DESI DR1 BAO with Pantheon+ Type Ia supernovae and Planck 2018 CMB likelihood components. Therefore, this section fits a Gaussian compression of the published joint posterior from that dataset combination.
4.2 External Posterior Chains and Gaussian Compression in $(w_{0},w_{a})$
@@
$\hat{\theta}\;=\;(\hat{w}_{0},\hat{w}_{a})^{T}$ (4.1)
and the weighted covariance matrix $\mathbf{C}$. We retain the off diagonal covariance and do not assume that $w_{0}$ and $w_{a}$ are independent. The DESI DR1 compressed likelihood products used in this work are publicly available. The analysis code script and derived output files are accessible as a .zip file at (Khan (2026)).
We obtain the following from running ancillary code script on cobaya MCMC chains,
$\hat{\theta}\;=\;\begin{pmatrix}\hat{w}_{0}\\ \hat{w}_{a}\end{pmatrix}\;=\;\begin{pmatrix}-0.828209\\ -0.744750\end{pmatrix},\qquad\mathbf{C}\;=\;\begin{pmatrix}0.004112&-0.016695\\ -0.016695&0.085227\end{pmatrix}$ (4.2)
4.3 Model Prediction $w_{\rm eff}(z)$ and mapping to an Effective CPL Pair